OpenStudy (hatahateo):

Help! I'll fan and medal Find the product of the complex number and its conjugate. 3 + 3i A. 18 B. 9 - 9i C. 0 D. 9 + 9i

OpenStudy (hatahateo):

@dumbcow @plainntall @welshfella @wkep1234

OpenStudy (plainntall):

So we have the complex number 3+3i what is its conjugate?

OpenStudy (hatahateo):

9-9i?

OpenStudy (plainntall):

No, the conjugate of a+bi is a-bi, you just change the sign of the imaginary part. So what is the conugate of 3+3i?

OpenStudy (hatahateo):

3-3i

OpenStudy (plainntall):

right, now we need to find the product of (3+3i)(3-3i)

OpenStudy (hatahateo):

18

OpenStudy (mww):

Simple rule for these If z is the complex number a + ib and z' is its conjugate a - ib Then their product is always a real number given by the square of the modulus! \[zz' = (a+ib)(a-ib) = a^2 - i^2b^2 = a^2 + b^2 = |z|^2\]

OpenStudy (plainntall):

Yes! :D

OpenStudy (hatahateo):

Thanks guys! Can you help me with one more? I'm a little confused, I'll open another question